OpenStudy (anonymous):

Describe each pattern using "math"/symbolsAND find the three next terms. 927,306,99,30,7,__,__,__

OpenStudy (anonymous):

There's not really a nice solution. It would look something like: \[ a_n=-\frac{1}{2}3^{2-n}(3^n-621) \]With the next three terms being: \[ -\frac{2}{3}, -\frac{29}{9}, -\frac{110}{27} \]